# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 4.2   Effective field theories
A central theme and relevant aspect for our purposes is the notion of composite hadrons and dynamically generated resonances. On physical grounds this refers to hadrons that are made of other hadrons, like the deuteron is a shallow bound state just below the proton-neutron threshold. In the meson sector, longstanding candidates for hadronic molecules are the light scalar mesons, the $\sigma/f_{0}(500)$, $\kappa/K^{*}_{0}(700)$, $a_{0}(980)$ and $f_{0}(980)$ [643]. Especially in the exotic heavy-meson sector there are now several well-established candidates for hadronic molecules [606]. On technical grounds, the realization of this phenomenon in EFTs is that ‘pre-existing’ hadrons, namely those appearing in the effective Lagrangian, can create new bound states or resonances by the strong channel couplings. For example, the light scalar mesons arise from the $\pi\pi$, $K\pi$ and $K\bar{K}$ couplings [632, 633, 631], and some states in the baryon sector from $N\pi$, $\Lambda K$, …interactions [636, 635]. In the light and strange baryon sector, the prime candidates for dynamically generated resonances are the $\Lambda(1405)$, the $N(1440)$ Roper resonance, and several other states [602, 603, 604, 605, 651]; for example, the authors of Refs. [636, 635] argue that also the $N(1535)$ and $N(1650)$ are dynamically generated.
@@
Sometimes it is argued that if a baryon resonance is dynamically generated it cannot be simultaneously a three-quark state, or at least its $qqq$ component should be small compared to $qqqq\bar{q}$ and higher multiquark components. Although this conclusion seems well justified in a number of cases, on general grounds one should advocate caution: Unless some quantum number prevents it, a baryon resonance must appear in any $qqq$, $qqqq\bar{q}$, …correlation function, so there is no dynamically generated baryon resonance without a three-quark core. Furthermore, one should also keep in mind that a BSE formulated in hadronic degrees of freedom does not know anything about quarks, so all it can do is create resonances dynamically, at least as long as the interactions allow it. For example, the $\pi\pi$, $K\bar{K}$, …interactions generate the light scalar mesons but at the same time also the $\rho$ and $K^{\ast}$, whose $q\bar{q}$ nature is undisputed. Unless one can somehow map the quark-gluon interactions in QCD to the hadronic interactions in EFTs, these are just two complementary toolboxes. Statements about compositeness require QCD calculations, e.g., like lattice QCD and functional methods where one can calculate the overlaps with $(qqq)$, $(qqq)(q\bar{q})$, …operators dynamically, or additional information such as the proximity to thresholds, decay patterns, line shapes, pole trajectories, large $N_{c}$ arguments, Regge phenomenology or unusual properties of elastic and transition form factors, see [643, 606, 604] and references therein.
Finally, a fruitful strategy which has gained traction over the last years is to combine EFTs with lattice QCD. Here the term ‘EFT’ does not necessarily refer to any specific effective theory derived from a Lagrangian, but rather to the general concept of employing hadronic scattering equations to satisfy unitarity. For example, this is needed to map finite-volume lattice data to infinite-volume scattering amplitudes in order to describe states above decay thresholds via the Lüscher method [652, 653]. Similarly, one can use BSEs to analytically continue lattice results for scattering amplitudes into the complex plane to determine the resonance pole locations. These aspects are further discussed in Sec. 4.6.
@@ 4.3   Quark models
Let us now adopt a different perspective and discuss baryon spectroscopy from microscopic approaches using quark and gluon degrees of freedom. The baseline for this are non-relativistic and relativistic constituent quark models, see [8, 560, 10, 654, 11, 530, 14, 1] for reviews. The idea is to write down a quantum-mechanical Hamiltonian for a system of $n$ valence quarks and/or antiquarks, in the non-relativistic version
$H=\sum_{i=1}^{n}\left(m_{i}+\frac{{\mbox{\boldmath$p$}}_{i}^{2}}{2m_{i}}\right)+\sum_{i<j}V({\mbox{\boldmath$r$}}_{ij})\,,$ (4.19)
[…]
Although we already mentioned some concrete models in Sec. 3, let us first discuss a few general features of the interquark potential $V({\mbox{\boldmath$r$}}_{ij})$. A useful starting point [655] is the Breit-Fermi interaction, i.e., the nonrelativistic expansion of the one-gluon exchange amplitude between quark and (anti)quark in QCD analogous to the one-photon exchange between electron and proton in the hydrogen atom:
$V({\mbox{\boldmath$r$}})=C\alpha_{s}\left[-\frac{1}{r}+V_{\text{ss}}({\mbox{\boldmath$r$}})+V_{\text{so}}({\mbox{\boldmath$r$}})+\dots\right].$ (4.20)
Besides the Coulomb and other spin-independent terms, it contains spin-dependent forces: The hyperfine potential $V_{\text{ss}}({\mbox{\boldmath$r$}})$ arises from the color-magnetic interaction and is responsible for the dominant mass splittings between states with different total quark spin $S$, e.g. between the pseudoscalar ($S=0$) and vector mesons ($S=1$) or the ground-state octet ($S=\nicefrac{{1}}{{2}}$) and decuplet ($S=\nicefrac{{3}}{{2}}$) baryons. The spin-orbit interaction $V_{\text{so}}({\mbox{\boldmath$r$}})$ does not contribute to orbital ground states ($L=0$) but gives rise to splittings for states with different $J$, e.g. for the charmonium states $\left\{\chi_{c0},\chi_{c1},\chi_{c2}\right\}$ with $J^{PC}=\{0,1,2\}^{++}$. By contrast, in the hydrogen atom the spin-orbit force beats the spin-spin interaction (due to $m_{e}\ll m_{p}$) and gives rise to the atomic fine structure effects.
@@
Quark confinement is usually implemented by adding a linearly rising potential $V_{\text{conf}}=\sigma r$ by hand to Eq. (4.20), where $\sigma$ is the string tension. This is motivated by Regge phenomenology and the observed mass spectra as well as lattice calculations of the Wilson loop for infinitely heavy quarks [656]. The dynamical origin of confinement is still under debate and several (not mutually exclusive) mechanisms have been proposed, e.g., center vortices in the path integral [657], ghost dominance in the infrared [658], or the formation of color-electric flux tubes and condensation of color-magnetic monopoles in the dual superconductor picture [659]. A hypothesis with interesting consequences for the electroweak sector is that confinement may be nothing more than gauge invariance [660]: Quarks and gluons depend on the gauge, but experimentally we can only measure gauge-invariant hadrons.
@@
Various quark models based on one-gluon exchange (OGE) interactions have been studied in the literature. The prototype model by de Rujula, Georgi and Glashow with spin-dependent forces reproduces the phenomenological Gell-Mann-Okubo relations for ground states [655]. A particularly influential model is the Isgur-Karl model [534, 539, 538, 536, 535, 537], which implements a confining harmonic oscillator potential together with an anharmonic perturbation and a hyperfine interaction. We already discussed the construction for baryons in Sec. 3.2: After removing the center-of-mass motion, one arrives at two independent spherical harmonic oscillators for the coordinates $\rho$ and $\lambda$. This leads to the band quantum number N and the pattern in Fig. 3.5 (right), where the mass splittings inside the $SU(6)$ multiplets are due to the hyperfine interaction and the splittings between the multiplets come from the anharmonic perturbation. The Isgur-Karl model describes the light and strange baryon spectrum reasonably well but also predicts more states than were observed at that time, which is the problem of the missing resonances. Capstick and Isgur used a relativized version of the model to fit the light and strange baryon spectrum with a consistent parameter set [548, 560]. Here the $N(1440)$ Roper resonance turns out to be heavier than the parity partner of the nucleon, the $N(1535)$. It was also found that many of the missing states couple only weakly to $N\pi$, which initiated photo- and electroproduction experiments around the world.
@@
The number of states in the spectrum can be reduced by assuming diquark clustering in baryons [661, 520, 662, 663, 521, 522]. For a pointlike diquark one of the oscillators in Eq. (3.21) is frozen, which reduces the number of excitations compared to the three-quark model. Because the total diquark wave function must be antisymmetric and its color part is already antisymmetric, for ground-state diquarks with a symmetric spatial wave function also the spin-flavor part must be symmetric. The spin-flavor $SU(6)$ diquark combination $\bm{6}\otimes\bm{6}=\bm{21}_{S}\oplus\bm{15}_{A}$ then only leaves the $\bm{21}$-plet, whose decomposition $\bm{21}={{}^{1}}\bar{\bm{3}}\oplus{{}^{3}}\bm{6}$ implies a scalar (‘good’) diquark with spin $S=0$ and an axialvector (‘bad’) diquark with $S=1$. Combining this with the remaining quark gives $\bm{21}\otimes\bm{6}=\bm{56}\oplus\bm{70}$, whereas the multiplets arising from $\bm{15}\otimes\bm{6}=\bm{20}\oplus\bm{70}$ are absent. On the other hand, since both $\bm{70}$-plets are needed to construct a totally antisymmetric baryon wave function, this also means that many states in the diquark model do not belong to irreducible representations of $SU(6)$; i.e., static diquarks that cannot exchange their roles also break the permutation symmetry. The simplest possibility of pointlike diquarks is also disfavored by comparison with experiment [411]. However, this does not prevent the formation of non-pointlike diquarks as two-body clusters that permanently exchange their roles through quark exchange, and there are indeed many indications for strong quark-quark correlations inside baryons [522].
@@
A different interaction mechanism is Goldstone-boson exchange (GBE) [664, 665], which emphasizes the role of chiral symmetry and the flavor dependence of the baryon spectrum. Here, the hyperfine interaction arises not from one-gluon exchange but from the exchange of octet pseudoscalar mesons between the quarks. This mechanism successfully predicts the correct level orderings of positive- and negative-parity nucleon excitations, such as the $N(1440)$ being lighter than the $N(1535)$. The relativistic version of the model provides a fair description of the spectrum and especially also various baryon form factors in the spacelike region [666, 667, 668, 592, 669, 670]. This is attributed to the importance of relativity, since in non-relativistic approaches form factors acquire relativistic recoil corrections at nonzero $Q^{2}$ which already affect their charge radii. A notable shortcoming here is the $\Lambda(1405)$, which comes out too high in mass, but this is a common problem among constituent-quark models.
@@
Figure 4.9: (Color online) Masses of nucleon (top) and $\Delta$ resonances (bottom) from quark model calculations. From left to right: Relativized one-gluon exchange by Capstick and Isgur [548], the algebraic model by Bijker, Iachello and Leviathan [565], the hypercentral quark model by Giannini, Santopinto and Vassallo [523], the relativistic Bonn model with instanton-induced interactions by Löring, Metsch and Petry [549], and its extension including Goldstone-boson exchange by Ronniger and Metsch [671]. The low-lying resonances in the first and second excitation shells are shown in darker colors and those in the third and fourth shells in lighter colors. The boxes give the experimental results for the Breit-Wigner masses with the uncertainty ranges from the RPP [3]. If no uncertainty estimate is given, or for new resonances, the experimental results from Refs. [672, 673, 474, 475] are shown.
@@
The hypercentral constituent quark model [674, 523, 675, 526, 528, 530] works with the hyperspherical coordinates $x=\sqrt{\rho^{2}+\lambda^{2}}$ and $\xi=\arctan\frac{\rho}{\lambda}$, where $\rho$ and $\lambda$ are the variables from Eq. (3.21), with $x$ representing the size of the state and $\xi$ its deformation. The potential is then expressed in these coordinates and the baryon wave functions are expanded in hyperspherical harmonics. The nonrelativistic version of the model with Coulomb, confinement, hyperfine and flavor-dependent interactions yields a good description of the spectrum and electromagnetic transition amplitudes [530, 528]. It also predicts more states than expected from experiment, and the flavor-dependent interaction is necessary to get the correct level ordering for the Roper resonance. The relativistic version of the model [526] also describes the elastic nucleon form factors well.
@@
Comprehensive calculations of the baryon spectrum have been achieved in the relativistic Bonn model employing instanton-induced interactions (I.I.I.) [676, 549, 677, 678]. Here one starts from the three-body Bethe-Salpeter or Faddeev equation discussed in Sec. 4.5 (see Fig. 4.13), except one employs free constituent quark propagators and the kernels are reduced to instantaneous interactions. This simplifies the structure of the equation to a Salpeter equation, which enables a similar treatment as in quark potential models. Quark confinement is implemented through a linearly rising three-quark string potential. Instead of the hyperfine interaction from one-gluon exchange, the authors employ a $qq$ interaction motivated by instanton effects, which accounts for the dominant mass splittings in the spectrum. A later extension [679, 671] also includes a flavor-dependent interaction similar to Goldstone-boson exchanges. With this setup, the light and strange baryon spectrum has been calculated up to high $J$ values. The model with flavor-dependent interactions correctly predicts the level ordering between positive- and negative-parity baryons, but the missing resonances problem remains. The model also gives a good description of various spacelike baryon elastic and transition form factors [568].
@@
Various other phenomenologically very successful models have been employed to describe the baryon excitation spectrum, including the ‘algebraic’ or collective models [565, 566, 680], where the residual interactions are described by a sum of spin-spin and spin-flavor operators, and large $N_{c}$ models [681, 682]. With the observations of exotic hadrons in the heavy-quark sector, quark models have also found widespread applications in the description of multiquark states in recent years, see e.g. [683, 684, 685, 686, 687, 688, 689, 690] for reviews.
@@
Fig. 4.9 shows the spectra of $N^{*}$ and $\Delta^{*}$ resonances calculated in different models in comparison to the experimental results. Instead of the real part of the pole position, we compare to the Breit-Wigner masses from the RPP [3] since the parameters of the calculations were chosen to reproduce those. The gross features of the low-lying states are well reproduced by all models. Below 1.7 GeV there are five negative-parity nucleon resonances and two $\Delta$ resonances, corresponding to the first shell in Fig. 3.2. In the positive-parity sector there is one low-mass state in the nucleon sector with $J^{P}=\nicefrac{{1}}{{2}}^{+}$ (the Roper) and one in the $\Delta$ sector with $J^{P}=\nicefrac{{3}}{{2}}^{+}$. Concerning the higher-lying positive parity states in the second shell, there are three states with $J^{P}=\nicefrac{{1}}{{2}}^{+}$, $\nicefrac{{3}}{{2}}^{+}$ and $\nicefrac{{5}}{{2}}^{+}$ around 1700 MeV in the nucleon sector and one $\Delta$ state, which may be identified with the (one-star) $\Delta(1750)\nicefrac{{1}}{{2}}^{+}$. In this region, the masses already vary considerably between the models. Above 1.9 GeV, numerous states are expected but only few have been found experimentally.
@@
The theoretical assumptions underlying these predictions are very different, and some deficiencies can be related to the underlying assumptions. The deficiencies can be seen inspecting the predicted masses of resonances in a few partial waves, see Table 4.1. For example, like many other models, the relativized OGE mechanism [548] has difficulties in explaining the low masses of the $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$. The mass difference between the Roper resonance and the negative-parity $N(1535)\nicefrac{{1}}{{2}}^{-}$ is $+90(34)$ MeV in experiment but $-60$ MeV in the calculation, and the difference between the $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ is +$140(73)$ MeV but $-175$ MeV in  [548]. The GBE model [666] with flavor-dependent interactions brings the mass of the Roper resonance below that of the $N(1535)\nicefrac{{1}}{{2}}^{-}$, however at the price of a somewhat mistuned fine structure: The mass separation between the $N(1650)\nicefrac{{1}}{{2}}^{-}$ and $N(1535)\nicefrac{{1}}{{2}}^{-}$ is $85(21)$ MeV compared to the calculated value of 254 MeV, while the difference between the $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ mass is $140(72)$ MeV but $-73$ MeV in the calculation. This implies that neither OGE nor GBE alone are sufficient to explain the spectrum.

### v4

@@ 4.2  Effective field theories
A central theme and relevant aspect for our purposes is the notion of composite hadrons and dynamically generated resonances. On physical grounds this refers to hadrons that are made of other hadrons, like the deuteron is a shallow bound state just below the proton-neutron threshold. In the meson sector, longstanding candidates for hadronic molecules are the light scalar mesons, the $\sigma/f_{0}(500)$, $\kappa/K^{*}_{0}(700)$, $a_{0}(980)$ and $f_{0}(980)$ Pelaez:2010fj . Especially in the exotic heavy-meson sector there are now several well-established candidates for hadronic molecules Guo:2017jvc . On technical grounds, the realization of this phenomenon in EFTs is that ‘pre-existing’ hadrons, namely those appearing in the effective Lagrangian, can create new bound states or resonances by the strong channel couplings. For example, the light scalar mesons arise from the $\pi\pi$, $K\pi$ and $K\bar{K}$ couplings Oller:1997ti , Oller:1998hw , Oller:1997ng , and some states in the baryon sector from $N\pi$, $\Lambda K$, …interactions Inoue:2001ip , Bruns:2010sv . In the light and strange baryon sector, the prime candidates for dynamically generated resonances are the $\Lambda(1405)$, the $N(1440)$ Roper resonance, and several other states Meissner:2020khl , Mai:2020ltx , Mai:2022eur , Guo:2023wes , Wang:2023snv . For example, the authors of Refs. Inoue:2001ip , Bruns:2010sv argue that also the $N(1535)$ and $N(1650)$ can be dynamically generated, see e.g. Li:2023pjx , Molina:2023jov for related discussions.
@@
Sometimes it is argued that if a baryon resonance is dynamically generated it cannot be simultaneously a three-quark state, or at least its qqqqqq component should be small compared to $qqqq\bar{q}$ and higher multiquark components. Although this conclusion seems well justified in a number of cases, on general grounds one should advocate caution: Unless some quantum number prevents it, a baryon resonance must appear in any qqqqqq, $qqqq\bar{q}$, …correlation function, so there is no dynamically generated baryon resonance without a three-quark core. Furthermore, one should also keep in mind that a BSE formulated in hadronic degrees of freedom does not know anything about quarks, so all it can do is create resonances dynamically, at least as long as the interactions allow it. For example, the $\pi\pi$, $K\bar{K}$, …interactions generate the light scalar mesons but at the same time also the $\rho$ and $K^{\ast}$, whose $q\bar{q}$ nature is undisputed. Unless one can somehow map the quark-gluon interactions in QCD to the hadronic interactions in EFTs, these are just two complementary toolboxes. Statements about compositeness require QCD calculations, e.g., like lattice QCD and functional methods where one can calculate the overlaps with $(qqq)$, $(qqq)(q\bar{q})$, …operators dynamically, or additional information such as the proximity to thresholds, decay patterns, line shapes, pole trajectories, large $N_{c}$ arguments, Regge phenomenology or unusual properties of elastic and transition form factors, see Pelaez:2010fj , Guo:2017jvc , Mai:2022eur and references therein.
Finally, a fruitful strategy which has gained traction over the last years is to combine EFTs with lattice QCD. Here the term ‘EFT’ does not necessarily refer to any specific effective theory derived from a Lagrangian, but rather to the general concept of employing hadronic scattering equations to satisfy unitarity. For example, this is needed to map finite-volume lattice data to infinite-volume scattering amplitudes in order to describe states above decay thresholds via the Lüscher method Luscher:1986pf , Luscher:1990ux . Similarly, one can use BSEs to analytically continue lattice results for scattering amplitudes into the complex plane to determine the resonance pole locations. These aspects are further discussed in Sec. 4.6.
@@ 4.3  Quark models
Let us now adopt a different perspective and discuss baryon spectroscopy from microscopic approaches using quark and gluon degrees of freedom. The baseline for this are non-relativistic and relativistic constituent quark models, see Hey:1982aj , Capstick:2000qj , Klempt:2009pi , Richard:2012xw , Crede:2013kia , Giannini:2015zia , Thiel:2022xtb , Gross:2022hyw for reviews. The idea is to write down a quantum-mechanical Hamiltonian for a system of $n$ valence quarks and/or antiquarks, in the non-relativistic version
$H=\sum_{i=1}^{n}\left(m_{i}+\frac{{\mbox{$p$}}_{i}^{2}}{2m_{i}}\right)+\sum_{i<j}V({\mbox{$r$}}_{ij})\,$, (4.19)
where the quarks carry constituent masses $m_{i}$ and the potential is the sum of pairwise interactions between the quarks with ${\mbox{$r$}}_{ij}={\mbox{$r$}}_{i}-{\mbox{$r$}}_{j}$. The solution of the Schrödinger equation $H\Psi_{\lambda}=E_{\lambda}\Psi_{\lambda}$ then determines the wave function $\Psi_{\lambda}$ of a hadron $|\lambda\rangle$ and its mass $E_{\lambda}=m_{\lambda}$ in the rest frame. The wave functions admit a direct probability interpretation through their modulus, and matrix elements are calculated via expectation values of operators with those wave functions.
Although we already mentioned some concrete models in Sec. 3, let us first discuss a few general features of the interquark potential $V({\mbox{$r$}}_{ij})$. A useful starting point DeRujula:1975qlm is the Breit-Fermi interaction, i.e., the nonrelativistic expansion of the one-gluon exchange amplitude between quark and (anti)quark in QCD analogous to the one-photon exchange between electron and proton in the hydrogen atom:
$V({\mbox{$r$}})=C\alpha_{s}\left[-\frac{1}{r}+V_{\text{ss}}({\mbox{$r$}})+V_{\text{so}}({\mbox{$r$}})+\dots\right].$ (4.20)
Besides the Coulomb and other spin-independent terms, it contains spin-dependent forces: The hyperfine potential $V_{\text{ss}}({\mbox{$r$}})$ arises from the color-magnetic interaction and is responsible for the dominant mass splittings between states with different total quark spin $S$, e.g. between the pseudoscalar ($S=0$) and vector mesons ($S=1$) or the ground-state octet ($S=\nicefrac{{1}}{{2}}$) and decuplet ($S=\nicefrac{{3}}{{2}}$) baryons. The spin-orbit interaction $V_{\text{so}}({\mbox{$r$}})$ does not contribute to orbital ground states ($L=0$) but gives rise to splittings for states with different $J$, e.g. for the charmonium states $\left\{\chi_{c0},\chi_{c1},\chi_{c2}\right\}$ with $J^{PC}=\{0,1,2\}^{++}$. By contrast, in the hydrogen atom the spin-orbit force beats the spin-spin interaction (due to $m_{e}\ll m_{p}$) and gives rise to the atomic fine structure effects.
@@
Quark confinement is usually implemented by adding a linearly rising potential $V_{\text{conf}}=\sigma r$ by hand to Eq. (4.20), where $\sigma$ is the string tension. This is motivated by Regge phenomenology and the observed mass spectra as well as lattice calculations of the Wilson loop for infinitely heavy quarks Bali:2000gf . The dynamical origin of confinement is still under debate and several (not mutually exclusive) mechanisms have been proposed, e.g., center vortices in the path integral Greensite:2003bk , ghost dominance in the infrared Alkofer:2000wg , or the formation of color-electric flux tubes and condensation of color-magnetic monopoles in the dual superconductor picture Kondo:2014sta . A hypothesis with interesting consequences for the electroweak sector is that confinement may be nothing more than gauge invariance Maas:2017wzi : Quarks and gluons depend on the gauge, but experimentally we can only measure gauge-invariant hadrons.
@@
Various quark models based on one-gluon exchange (OGE) interactions have been studied in the literature. The prototype model by de Rujula, Georgi and Glashow with spin-dependent forces reproduces the phenomenological Gell-Mann-Okubo relations for ground states DeRujula:1975qlm . A particularly influential model is the Isgur-Karl model Isgur:1977ef , Isgur:1979be , Isgur:1978xb , Isgur:1978xi , Isgur:1978xj , Isgur:1978wd , which implements a confining harmonic oscillator potential together with an anharmonic perturbation and a hyperfine interaction. We already discussed the construction for baryons in Sec. 3.2: After removing the center-of-mass motion, one arrives at two independent spherical harmonic oscillators for the coordinates $\rho$ and $\lambda$. This leads to the band quantum number N and the pattern in Fig. 3.5 (right), where the mass splittings inside the $SU(6)$ multiplets are due to the hyperfine interaction and the splittings between the multiplets come from the anharmonic perturbation. The Isgur-Karl model describes the light and strange baryon spectrum reasonably well but also predicts more states than were observed at that time, which is the problem of the missing resonances. Capstick and Isgur used a relativized version of the model to fit the light and strange baryon spectrum with a consistent parameter set Capstick:1986ter , Capstick:2000qj . Here the $N(1440)$ Roper resonance turns out to be heavier than the parity partner of the nucleon, the $N(1535)$. It was also found that many of the missing states couple only weakly to $N\pi$, which initiated photo- and electroproduction experiments around the world.
@@
The number of states in the spectrum can be reduced by assuming diquark clustering in baryons Lichtenberg:1969sxc , Anselmino:1992vg , Santopinto:2004hw , Ebert:2007nw , Santopinto:2014opa , Barabanov:2020jvn . For a pointlike diquark one of the oscillators in Eq. (3.21) is frozen, which reduces the number of excitations compared to the three-quark model. Because the total diquark wave function must be antisymmetric and its color part is already antisymmetric, for ground-state diquarks with a symmetric spatial wave function also the spin-flavor part must be symmetric. The spin-flavor $SU(6)$ diquark combination $\bm{6}\otimes\bm{6}=\bm{21}_{S}\oplus\bm{15}_{A}$ then only leaves the $\bm{21}$-plet, whose decomposition $\bm{21}={{}^{1}}\bar{\bm{3}}\oplus{{}^{3}}\bm{6}$ implies a scalar (‘good’) diquark with spin $S=0$ and an axialvector (‘bad’) diquark with $S=1$. Combining this with the remaining quark gives $\bm{21}\otimes\bm{6}=\bm{56}\oplus\bm{70}$, whereas the multiplets arising from $\bm{15}\otimes\bm{6}=\bm{20}\oplus\bm{70}$ are absent. On the other hand, since both $\bm{70}$-plets are needed to construct a totally antisymmetric baryon wave function, this also means that many states in the diquark model do not belong to irreducible representations of $SU(6)$; i.e., static diquarks that cannot exchange their roles also break the permutation symmetry. The simplest possibility of pointlike diquarks is also disfavored by comparison with experiment Nikonov:2007br . However, this does not prevent the formation of non-pointlike diquarks as two-body clusters that permanently exchange their roles through quark exchange, and there are indeed many indications for strong quark-quark correlations inside baryons Barabanov:2020jvn .
@@
A different interaction mechanism is Goldstone-boson exchange (GBE) Glozman:1995fu , Glozman:1996wq , which emphasizes the role of chiral symmetry and the flavor dependence of the baryon spectrum. Here, the hyperfine interaction arises not from one-gluon exchange but from the exchange of octet pseudoscalar mesons between the quarks. This mechanism successfully predicts the correct level orderings of positive- and negative-parity nucleon excitations, such as the $N(1440)$ being lighter than the $N(1535)$. The relativistic version of the model provides a fair description of the spectrum and especially also various baryon form factors in the spacelike region Glozman:1997ag , Glozman:2001zc , Melde:2008yr , Melde:2007zz , Melde:2008dg , Plessas:2015mpa . This is attributed to the importance of relativity, since in non-relativistic approaches form factors acquire relativistic recoil corrections at nonzero $Q^{2}$ which already affect their charge radii. A notable shortcoming here is the $\Lambda(1405)$, which comes out too high in mass, but this is a common problem among constituent-quark models.
@@
Figure 4.9: (Color online) Masses of nucleon (top) and $\Delta$ resonances (bottom) from quark model calculations. From left to right: Relativized one-gluon exchange by Capstick and Isgur Capstick:1986ter , the algebraic model by Bijker, Iachello and Leviathan Bijker:1994yr , the hypercentral quark model by Giannini, Santopinto and Vassallo Giannini:2001kb , the relativistic Bonn model with instanton-induced interactions by Löring, Metsch and Petry Loring:2001kx , and its extension including Goldstone-boson exchange by Ronniger and Metsch Ronniger:2012xp . The low-lying resonances in the first and second excitation shells are shown in darker colors and those in the third and fourth shells in lighter colors. The boxes give the experimental results for the Breit-Wigner masses with the uncertainty ranges from the RPP ParticleDataGroup:2024cfk . If no uncertainty estimate is given, or for new resonances, the experimental results from Refs. BES:2009ufh , BESIII:2012ssm , CLAS:2024iir , Sarantsev:2025lik are shown.
@@
The hypercentral constituent quark model Ferraris:1995ui , Giannini:2001kb , Giannini:2005ks , DeSanctis:2005kt , Santopinto:2012nq , Giannini:2015zia works with the hyperspherical coordinates $x=\sqrt{\rho^{2}+\lambda^{2}}$ and $\xi=\arctan\frac{\rho}{\lambda}$, where $\rho$ and $\lambda$ are the variables from Eq. (3.21), with $x$ representing the size of the state and $\xi$ its deformation. The potential is then expressed in these coordinates and the baryon wave functions are expanded in hyperspherical harmonics. The nonrelativistic version of the model with Coulomb, confinement, hyperfine and flavor-dependent interactions yields a good description of the spectrum and electromagnetic transition amplitudes Giannini:2015zia , Santopinto:2012nq . It also predicts more states than expected from experiment, and the flavor-dependent interaction is necessary to get the correct level ordering for the Roper resonance. The relativistic version of the model DeSanctis:2005kt also describes the elastic nucleon form factors well.
@@
Comprehensive calculations of the baryon spectrum have been achieved in the relativistic Bonn model employing instanton-induced interactions (I.I.I.) Loring:2001kv , Loring:2001kx , Loring:2001ky , Metscha:2008gkf . Here one starts from the three-body Bethe-Salpeter or Faddeev equation discussed in Sec. 4.5 (see Fig. 4.13), except one employs free constituent quark propagators and the kernels are reduced to instantaneous interactions. This simplifies the structure of the equation to a Salpeter equation, which enables a similar treatment as in quark potential models. Quark confinement is implemented through a linearly rising three-quark string potential. Instead of the hyperfine interaction from one-gluon exchange, the authors employ a $qq$ interaction motivated by instanton effects, which accounts for the dominant mass splittings in the spectrum. A later extension Ronniger:2011td , Ronniger:2012xp also includes a flavor-dependent interaction similar to Goldstone-boson exchanges. With this setup, the light and strange baryon spectrum has been calculated up to high $J$ values. The model with flavor-dependent interactions correctly predicts the level ordering between positive- and negative-parity baryons, but the missing resonances problem remains. The model also gives a good description of various spacelike baryon elastic and transition form factors Merten:2002nz .
@@
Various other phenomenologically very successful models have been employed to describe the baryon excitation spectrum, including the ‘algebraic’ or collective models Bijker:1994yr , Bijker:1995ii , Bijker:2000gq , where the residual interactions are described by a sum of spin-spin and spin-flavor operators, and large $N_{c}$ models Goity:2003ab , Matagne:2004pm . With the observations of exotic hadrons in the heavy-quark sector, quark models have also found widespread applications in the description of multiquark states in recent years, see e.g. Brambilla:2019esw , Liu:2019zoy , Dong:2021bvy , Dong:2021juy , Chen:2022asf , Burns:2022uiv , Klempt:2022zwo , Liu:2024uxn for reviews.
@@
Fig. 4.9 shows the spectra of $N^{*}$ and $\Delta^{*}$ resonances calculated in different models in comparison to the experimental results. Instead of the real part of the pole position, we compare to the Breit-Wigner masses from the RPP ParticleDataGroup:2024cfk since the parameters of the calculations were chosen to reproduce those. The gross features of the low-lying states are well reproduced by all models. Below 1.7 GeV there are five negative-parity nucleon resonances and two $\Delta$ resonances, corresponding to the first shell in Fig. 3.2. In the positive-parity sector there is one low-mass state in the nucleon sector with $J^{P}=\nicefrac{{1}}{{2}}^{+}$ (the Roper) and one in the $\Delta$ sector with $J^{P}=\nicefrac{{3}}{{2}}^{+}$. Concerning the higher-lying positive parity states in the second shell, there are three states with $J^{P}=\nicefrac{{1}}{{2}}^{+}$, $\nicefrac{{3}}{{2}}^{+}$ and $\nicefrac{{5}}{{2}}^{+}$ around 1700 MeV in the nucleon sector and one $\Delta$ state, which may be identified with the (one-star) $\Delta(1750)\nicefrac{{1}}{{2}}^{+}$. In this region, the masses already vary considerably between the models. Above 1.9 GeV, numerous states are expected but only few have been found experimentally.
@@
The theoretical assumptions underlying these predictions are very different, and some deficiencies can be related to the underlying assumptions. The deficiencies can be seen inspecting the predicted masses of resonances in a few partial waves, see Table 4.3. For example, like many other models, the relativized OGE mechanism Capstick:1986ter has difficulties in explaining the low masses of the $N(1440)\nicefrac{{1}}{{2}}^{+}$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$. The mass difference between the Roper resonance and the negative-parity $N(1535)\nicefrac{{1}}{{2}}^{-}$ is +90(34)+90(34) MeV in experiment but 60-60 MeV in the calculation, and the difference between the $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ is +$140(73)$ MeV but 175-175 MeV in  Capstick:1986ter . The GBE model Glozman:1997ag with flavor-dependent interactions brings the mass of the Roper resonance below that of the $N(1535)\nicefrac{{1}}{{2}}^{-}$, however at the price of a somewhat mistuned fine structure: The mass separation between the $N(1650)\nicefrac{{1}}{{2}}^{-}$ and $N(1535)\nicefrac{{1}}{{2}}^{-}$ is $85(21)$ MeV compared to the calculated value of 254 MeV, while the difference between the $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$ and $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ mass is $140(72)$ MeV but 73-73 MeV in the calculation. This implies that neither OGE nor GBE alone are sufficient to explain the spectrum.